Shift report 20010920 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Thu Sep 20 2001 - 15:59:46 EDT

  • Next message: Flemming Videbaek: "shitfs --- HV problems"

    ------------------------------------------------------------
    Date: Thu Sep 20 15:59:45 2001, Shift: 08:00-16:00
    Supervisor: Eun-Joo Kim, with: Burton Budick
    Log book#9 pages 14 - 16
    
    Shift Summary:
    ------------------------------------------------------------
    1. No beam during entire shift
    
    2. Run 4823 : data run with 90k trigger 1
       Run 4824, 4825 : Pedestal run for camac, TPC
       Run 4827 : DVM run
    
    3. Mainframe tile4, tile5 have problems.
       Bob and Ramiro are working on HV problems.
    
       Ramiro refilled RICH, pressure is 970.0 Torr.
       Hiro made a macro for DVM run.
       
       Magnet technican checked power supply for D5 around 2pm.
       There was no problem to reading magnet value back.   
    
       
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Thu Sep 20 2001 - 16:00:32 EDT